OpenStudy (anonymous):

Please help me do this homework problem, I don't understand maths at all so I have been stuck on it for two days and if I submit it blank my grade will be an F :( Find all the zeroes of the equation. x^4–6x^2–7x–6=0

OpenStudy (anonymous):

Let f(x) = x^4 – 6x^2 – 7x – 6 Note factors of the constant term –6 are ±1, ±2, ±3 and ±6 f(−2) = 0 means x+2 is a factor of f(x) f(3) = 0 means x−3 is a factor of f(x) Do long division or synthetic division or by inspection f(x) = (x+2)(x−3)(x²+x+1) (x+2)(x−3)(x²+x+1) = 0 x = −2 or x = 3 or x²+x+1 = 0 Solving the quadratic equation, in addition to the 2 real solutions, x = ½(−1 ± i√3).

OpenStudy (anonymous):

did you copy this from yahoo answers? http://answers.yahoo.com/question/index?qid=20130429222954AAogGfU

OpenStudy (anonymous):

I asked here because I didn't fully understand the explanation on yahoo answers

OpenStudy (anonymous):

idk how to explain it. Try asking your teacher.

OpenStudy (luigi0210):

Any specific method?

OpenStudy (anonymous):

@Luigi0210 no specific method that i know of. the assignment just says do it and show my work (but the only way i know how to do is use a calculator)

OpenStudy (anonymous):

im just going to do it with a calculator. i will only get 1 point taken off and shouldn't drop to an F

OpenStudy (luigi0210):

Yea, just graph it and locate points of intersection.. and sorry, OS is still messed up so i got no notification. ლ(ಠ_ಠლ)